OpenStudy (anonymous):

y^2 + 3yz - 8z -4x = 0 Solve for z...

OpenStudy (anonymous):

y^2 + 3yz - 8z - x = 0 Isolate the z's. 3yz - 8z = x - y^2 Factor out z. z(3y - 8) = x - y^2 Divide by 3y - 8. z = (x - y^2) / (3y - 8)

OpenStudy (anonymous):

thank you!